# Evaluation of uncertainty estimations for Gaussian process regression based machine learning interatomic potentials

### v1

@@ Abstract
Machine learning interatomic potentials (MLIPs) have seen significant advances as efficient replacement of expensive quantum chemical calculations for predicting molecular energies and forces. Uncertainty estimations for these machine learning models are crucial to quantify the additional model error they introduce and to leverage this information to find most informative training samples in active learning strategies. MLIPs that are based on Gaussian process regression provide predictive distributions with a standard deviation as a possible uncertainty measure. An alternative approach are ensemble-based uncertainties. Although these uncertainty measures have been applied to active learning, it has rarely been studied how well or whether they correlate with the actual error, and it is not always clear whether active learning strategies actually outperform randomized sampling strategies. In our study, we consider GPR models with Coulomb and Smooth Overlap of Atomic Positions (SOAP) representations as inputs to predict potential energy surfaces and excitation energies of molecules. We evaluate, how the GPR variance and ensemble-based uncertainties relate to the actual error and whether model performance improves by selecting the most uncertain samples from a fixed configuration space for training. For the ensemble based uncertainty estimations, we find that they often do not provide any information about the error. For the predictive GPR standard deviation, we find that in most cases predictions with an increasing standard deviation also have an increasing systematical bias, which is not captured by the uncertainty. In these cases, selecting training samples with the highest uncertainty leads to a model with a worse test error compared to selecting samples randomly. We conclude that confidence intervals for GPR-based MLIPs, which are derived from the predictive standard deviation, can be highly overconfident, as they do not capture the bias in the model predictions. Selecting samples with high GPR standard deviation leads to a model that overemphasizes the borders of the configuration space represented in the fixed dataset. This may result in worse performance in more densely sampled areas but better generalization for extrapolation tasks.
@@ 1 Introduction
How well samples from different parts of the configuration space are predicted by an MLIP depends highly on its training data. By configuration space, we refer to the set of all possible spatial arrangements of atoms in an atomic system. Samples from areas of the configuration space that differ too much from those covered by the training data might not be predicted well. Because of this variability in prediction accuracy, it becomes desirable to provide an uncertainty estimation for every model prediction, which ideally gives a good estimation for the prediction error. A good error estimation could help to understand the range of a potential error of a prediction, and could be utilized in active learning (AL) strategies1; 2; 3; 4; 5; 6; 7; 8; 9; 10; 11; 12; 13; 14; 15; 16; 17. With AL, algorithms are referred to that aim to choose or create new training samples, which are maximally informative to the model. In the context of MLIPs, two important goals of AL exist. The first goal is to increase the data efficiency of the model training. Data efficiency refers to the reduction in the number of training samples needed to achieve comparable model accuracy within a fixed configuration space, compared to using randomly selected samples. A second goal is to sample the configuration space, without being fixed, in all areas which are relevant for the respective applications.
@@
Many different ML approaches to design MLIPs exist. One class of MLIPs are message passing neural networks 18; 19; 20; 21 (MPNN), which contain features that are trained via message passing steps. In other approaches, instead of trained features, fixed representations are employed that either globally represent the whole system of atoms 22; 23; 24 or locally represent atomistic environments 25; 26; 27; 28. These representations provide important properties such as invariance against translation, rotation, and atom permutation29. Besides linear models 26; 30 and neural networks28; 31; 32, kernel-based models are commonly chosen as ML model 33; 22; 24. Within kernel-based models, many are variations of Gaussian process regression34 (GPR). GPR predicts the output for a new input by generating a Gaussian predictive distribution, where the mean represents the (deterministic) prediction and the variance provides a measure of uncertainty for the prediction. The GPR variance is commonly used as an uncertainty measure in AL schemes for GPR-based MLIPs9; 16; 3; 1. Besides the GPR variance, other uncertainty estimations are conceivable for GPR. A common approach in ML is to construct ensembles of slightly different models and use the variance of their predictions as an uncertainty measure. Model ensembles can be based on different initializations of weights 35 or they can be generated by training on different subsets of the training data36. For neural networks, using different weight initializations is reasonable because these models are parametric and sensitive to their initial conditions. In contrast, for GPR, it is more reasonable to use different subsets of the training data because GPR is a non-parametric method that relies directly on the training data to make predictions. Uteva et al.9 trained two models on half of the initial training set each. Similarly, a bootstrap aggregation36; 37 ensemble, which is commonly used in ML, would be possible for GPR. Here multiple models are trained on different subsets of the training data, which are generated by randomly sampling indices from the original dataset. Our focus will be on GPR-based MLIPs with GPR variance and ensemble approaches as uncertainty measure or predictive distributions.
GPR models, as well as GPR ensembles, provide predictive distributions with a mean and a standard deviation. The mean of this distribution may serve as point prediction, which is the model’s actual predicted value. The standard deviation is a measure of variability or uncertainty in the prediction. An important question is, to which extend such uncertainty measures are correlated to the prediction error of a model, thus whether they would serve as error estimator. This requires a careful evaluation. Evaluating uncertainty estimations of MLIPs is a challenging problem with a large variety of approached being reported in the literature. Tran et al.38 distinguished between three aspects of uncertainty estimation: calibration, sharpness and dispersion. Calibration describes if the uncertainty estimation is correct, in the sense that the true error values follow the predicted distribution. Sharpness characterizes the overall confidence of a model’s predictions, and generally, a model with more confident predictions is preferred. Dispersion, on the other hand, refers to the spread of the uncertainty estimates. A model with high dispersion shows a wide variability in confidence across different predictions, allowing it to differentiate between samples predicted with high accuracy and those potentially having high errors. Tran et al.38 were analyzing these aspects with a variety of metrics. In other studies, often the uncertainty estimations are plotted against the actual or absolute errors3; 39; 40; 41; 10, which, in principal, gives an intuitively accessible picture off all three aspects. Kellner et al. 39 constructed an example with synthetic data to show how an ideal uncertainty versus error double-logarithmic plot would look like and illustrated various cases of incorrect uncertainty estimations. They then analyzed, if uncertainty versus error plots for non-synthetic data show the correct shape. Busk et al. 40 sorted all predictions in equally sized bins and calculated metrics like the z-score variance and the RMSE for all samples of one bin. In this work, we use binning to analyze the error distributions of predictions with certain uncertainties.
In the literature, uncertainty estimations of MLIPs are applied in a large variety of different active learning schemes. Many of these create new configurations without explicitly restricting the configuration space. A particularly important AL scheme is on-the-fly AL. Here, an uncertainty is calculated at every point during an MD run with the MLIP to recognize if the MLIP runs into configuration areas which are poorly sampled in the training data and for which the MD run might become instable1; 2; 3; 4; 5; 6; 7. Other schemes are optimizing the uncertainty estimation multiplied by a weight function to create new training configurations15; 16; 17. However, when using uncertainty estimation to guide the exploration of a flexible configuration space, it becomes challenging to assess the actual impact of active learning and the chosen uncertainty metric in comparison to generating samples randomly. Therefore, we restrict ourselves to a simple, very common AL approach from standard ML literature, which samples a fixed configuration space. This approach is often referred to as uncertainty sampling 42; 9. Uncertainty sampling can be applied in setups, where a large amount of unlabeled data exists and generating the respective labels is expensive. For a small random subset of the data, labels are generated. Then this data is used to train an initial model. The uncertainty estimation of this model is used to iteratively select samples from the remaining pool of unlabeled samples for which the labels are generated and which are then added as additional training samples. In the general ML literature, uncertainty sampling is often applied to classification problems, but it is also used for regression problems 43. It was shown that uncertainty sampling can increase the data efficiency by a multiple, although the effect strongly depends on the dataset and the uncertainty estimation that is used. In this study, uncertainty sampling is applied to analyze the benefit of the discussed different uncertainty estimation methods on AL.
In literature, uncertainty estimations of GPR based MLIPs are often evaluated by directly examining the performance when applied to AL without explicitly analyzing how the uncertainty estimations correlate with the actual error9; 1; 16. These studies give a mixed picture on the effect of AL on training GPR-based MLIPs and it is often not clear, whether AL strategies provide an actual improvement over random sampling. The correlation between uncertainty estimations of GPR-based MLIPs and the actual error has rarely been studied, explicitly. Vandermause et al.3 plotted their Gaussian process uncertainty against the true absolute error and reported that most true absolute errors fall within twice the Gaussian process uncertainty. They further showed how their active learning method, based on the Gaussian process uncertainty, provides actual improvement over random sampling. Bartok et al. 44 plotted uncertainty estimations against the true absolute error for a dataset of Ar dimers and trimers, finding that maximizing the leave-one-out cross-validation likelihood for hyperparameter optimization can be advantageous for uncertainty estimation, compared to maximizing the marginal likelihood or manual hyperparameter tuning. Tran et al. 38 evaluated different aspects of uncertainty estimations for an exemplary dataset using, among others, a GPR-based MLIP. Various GPR-based MLIPs exist, which differ in how their kernel is constructed and which representations are used. Therefore, also the GPR variance differs and it is not clear how the GPR variances of different models compare as uncertainty estimations. Furthermore, even if an uncertainty estimation shows a meaningful correlation with the errors, translating this relationship into enhanced data efficiency for AL runs remains uncertain.
The objective of this study is hence to evaluate uncertainty measures for GPR-based MLIPs. We consider a GPR model with global Coulomb representations22 as inputs as well as an atomistic GPR model with SOAP25 representations. We test the prediction of ground state energies from molecules of well-established benchmarks, allowing for easy comparison and categorization of our results. We further predict excited state energies for a trajectory of porphyrin in gas phase as a large and complex molecule with a potentially complex configuration space. To evaluate the uncertainty estimations, we test how the GPR variance, as well as ensemble uncertainties, of different GPR-based MLIPs, correlate with the actual error for unseen samples. This is done by analyzing the distributions of errors of predictions with different uncertainties. It is further investigated how different uncertainty measures perform, when applied to AL. Uncertainty sampling is chosen as AL algorithm for the reasons discussed above. To understand, how much information could be maximally extracted from the candidate pool, when using uncertainty sampling, we always also provide a reference AL experiment. In that reference AL experiment, exact information on the error is used, in contrast to using an uncertainty measure.
This work is structured as follows. In Section 2, Gaussian Process Regression (GPR) is introduced, and the use of Coulomb and SOAP representations to construct GPR-based MLIPs is described. The different uncertainty estimations that are applied are described and we give a detailed description of how we evaluate them. Further details are given on how uncertainty sampling was applied and which data is used. In Section 3 we analyze the correlation of uncertainty estimations and errors for the different datasets. It is further investigated how the observed correlations transfer when applied to uncertainty sampling. An overall discussion of the results is given in Section 4. Conclusions with an outlook are given in Section 5.
@@ 2.1 Gaussian process regression
In GPR34, a prior probability distribution over functions is defined using a Gaussian process. We want to find a posterior predictive distribution from training inputs $\mathbf{x}_{i}$ and corresponding output measurements $y_{i}$, with which we can predict outputs for unknown inputs x. Our output measurements will always be noisy. This can be considered by assuming that the actual function values differ from the measurements with
@@
where $\epsilon_{i}$ are independent, identically distributed random variables drawn from a Gaussian distribution with zero mean and variance $\sigma_{n}^{2}$. The GPR framework provides a predictive distribution for the output corresponding to any input x. The point prediction for the output of an input $\mathbf{x}$ is the mean of this predictive distribution, given as
@@
which is one of the most common choices. We used a different length-scale for every feature, known as automatic relevance determination34, to increase the accuracy and flexibility of the model. Here $\mathbf{M}$ is a diagonal matrix $\text{diag}\left(\frac{1}{l_{1}^{2}},...,\frac{1}{l_{D}^{2}}\right)$, with $D$ the number of features, and $\sigma_{f}^{2}$ is the output-scale. We optimized the hyperparameters $\sigma_{f}^{2}$ and $\sigma_{n}^{2}$, as well as length-scales inside the covariance function via maximum marginal log-likelihood, details can be found in Section S2. For all our tests, we used the GPR implementation of GPyTorch 45.
In the case of a global representation as inputs $\mathbf{x}$, the covariance function of Eq. 3 is directly applied to these inputs. For our study, we incorporate Coulomb matrices22 as a global representation. For Coulomb matrices, to ensure invariance against atom indexing, often a sorting is carried out, which results in discontinuities, which may cause challenges in the construction of ML models 29. In our tests, we ignored invariance against atom permutation and omitted sorting because, for all tasks, a system with fixed atoms given in a fixed order was given.
As a local representation, we used Smooth overlap of atomic positions (SOAP) 25. It is commonly used and GPR models with SOAP representations as inputs have shown high accuracy 29. For atomistic representations, an input x will consist of representations $\textbf{x}_{j}$ for every atomistic environment of the system. The covariance function between two inputs x and x’ is built by summing over an atomistic covariance function
@@ 2.2 Uncertainty estimations
with a given mean $m(\mathbf{x})$, and a standard deviation $u(\mathbf{x})$, which we consider as uncertainty measure in here. Note that we may interpret the pair from eq. 5 as the parameters describing a Gaussian distribution. In fact, the predictive distribution of a GPR model is anyway Gaussian. Moreover, the predictive distributions of GPR ensembles are often assumed to be Gaussian too40; 47, which is, according to the central limit theorem, at least true for a large number of ensemble models47.
@@
where $m_{g}(\mathbf{x})$ is the GPR mean from eq. 2 and we have
$u_{g}\left(\textbf{x}\right)=\sigma\left(\textbf{x}\right)$ (7)
for all $\mathbf{x}$, with the GPR variance
$\sigma^{2}\left(\mathbf{x}\right)=k\left(\mathbf{x},\mathbf{x}\right)-\mathbf{K}\left(\mathbf{x},\mathbf{X}\right)\bigl{(}\mathbf{K}\left(\mathbf{X},\mathbf{X}\right)+\sigma_{n}^{2}\mathbf{I}\bigr{)}^{-1}\mathbf{K}\left(\mathbf{X},\mathbf{x}\right)\,.$ (8)
Note that the GPR variance is only calculated from the kernel function $k(\cdot,\cdot)$, the inputs $\mathbf{X}$ of the training set, i.e. the sample locations in the feature space, and the noise variance $\sigma_{n}^{2}$. Hence it does not depend on the training outputs, i.e. the trained energies. Intuitively, the GPR variance is high, when $\mathbf{x}$ is far from the training inputs, as measured by the GPR model. In the variance, see eq. 8, GPR already accounts for the noise $\sigma_{n}^{2}$ in the observations, therefore, we do not add this noise to our uncertainty estimation, like it was done by Vandermause et al.3.
The two set approach, proposed by Uteva et al.9, was motivated by the principles underlying the jackknife and bootstrap methods36. We use for its predictive distribution the notation
@@
is the uncertainty estimation. The square of this particular uncertainty measure can be interpreted as a very rough estimate of the model’s variance with respect to changes in the training set, including the energies. In other words, it very roughly approximates the variance in the bias-variance decomposition of ML models, which is common in ML literature48.
An extension of this is the bootstrap aggregation ensemble36, which leads to the predictive distribution
@@
${m}_{b}\left(\textbf{x}\right)=\sum_{l=1}^{N}m^{\left(l\right)}\left(\textbf{x}\right)\,,$ (13)
@@
As a slight change towards the classical definition of the two-sets and bootstrap approach, we do not use the ensemble mean as mean for their predictive distributions, since we found whether we use the mean of the ”full” GPR model or the ensemble mean does not make a difference for the essential results on the correlation between error and uncertainty estimation. A comparison can be found in Section S2.

### v2

@@ Abstract
Uncertainty estimations for machine learning interatomic potentials (MLIPs) are crucial for quantifying model error and identifying informative training samples in active learning strategies. In this study, we evaluate uncertainty estimations of Gaussian process regression (GPR)-based MLIPs, including the predictive GPR standard deviation and ensemble-based uncertainties. We do this in terms of calibration and in terms of impact on model performance in an active learning scheme. We consider GPR models with Coulomb and Smooth Overlap of Atomic Positions (SOAP) representations as inputs to predict potential energy surfaces and excitation energies of molecules. Regarding calibration, we find that ensemble-based uncertainty estimations show already poor global calibration (e.g., averaged over the whole test set). In contrast, the GPR standard deviation shows good global calibration, but when grouping predictions by their uncertainty, we observe a systematical bias for predictions with high uncertainty. Although an increasing uncertainty correlates with an increasing bias, the bias is not captured quantitatively by the uncertainty. Therefore, the GPR standard deviation can be useful to identify predictions with a high bias and error but, without further knowledge, should not be interpreted as a quantitative measure for a potential error range. Selecting the samples with the highest GPR standard deviation from a fixed configuration space leads to a model that overemphasizes the borders of the configuration space represented in the fixed dataset. This may result in worse performance in more densely sampled areas but better generalization for extrapolation tasks.
@@ 1 Introduction
How well samples from different areas of the configuration space are predicted by an MLIP depends highly on its training data. Samples from areas of the configuration space (e.g., spatial arrangements of atoms in an atomic system) that differ too much from those covered by the training data might not be predicted well. Because of this variability in prediction accuracy, it becomes desirable to provide an uncertainty estimation for every model prediction. A good uncertainty estimation could help to understand the range of a potential error of a prediction, and be utilized in active learning (AL) strategies[1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17]. With AL, algorithms are referred to that aim to choose or create new training samples, which are maximally informative to the model.
@@
Many different ML approaches to design MLIPs exist. One class of MLIPs are message passing neural networks [18, 19, 20, 21] (MPNN), which contain features that are trained via message passing steps. In other approaches, instead of trained features, fixed representations are employed that either globally represent the whole system of atoms [22, 23, 24] or locally represent atomistic environments [25, 26, 27, 28]. These representations provide important properties such as invariance against translation, rotation, and atom permutation[29]. Besides linear models [26, 30] and neural networks[28, 31, 32], kernel-based models are commonly chosen as ML model [33, 22, 24]. Within kernel-based models, many are variations of Gaussian process regression[34] (GPR). GPR predicts the output for a new input by generating a Gaussian predictive distribution, where the mean represents the (deterministic) prediction, and the standard deviation provides a measure of uncertainty for the prediction. The GPR standard deviation is commonly used as an uncertainty measure in AL schemes for GPR-based MLIPs[9, 16, 3, 1]. Besides the GPR standard deviation, other uncertainty estimations are conceivable for GPR. A common approach in ML is to construct ensembles of slightly different models and use the standard deviation of their predictions as an uncertainty measure. Since GPR is a non-parametric method that relies directly on the training data for predictions, it is reasonable to construct ensembles where the models differ based on subsets of the training data. Uteva et al.[9] trained two models on half of the initial training set each. Similarly, a bootstrap aggregation[35, 36] ensemble, which is commonly used in ML, can be applied to GPR.
Both the GPR standard deviation and ensemble-based uncertainties are standard deviations of predictive distributions, thus provide uncertainty estimations. To evaluate these uncertainty estimations, the most important aspect is calibration, which describes how well the distribution of empirical errors aligns with the predictive distributions. A common tool to evaluate calibration are calibration curves[37], which test whether the empirical errors are covered by the confidence intervals derived from the predictive distributions. However, it has been pointed out that this approach does not sufficiently validate good calibration, because it averages over the entire test dataset [38, 39]. Separating predictions into bins by their uncertainty can give a more localized evaluation of uncertainty estimations, for example, using reliability diagrams[39], or by calculating the variance of z-scores[38]. In the literature evaluating uncertainty estimations of general MLIPs, local analyses of calibration are often lacking or limited to qualitative tests, such as plotting errors against uncertainties [40, 41, 10, 42]. In contrast, Busk et al. [43] sorted all predictions into bins to generate reliability diagrams and calculate local z-score variances. In this work, we use calibration curves as well as extended reliability diagrams to evaluate calibration. While classical reliability diagrams only show the standard deviation of empirical errors within each bin, extended reliability diagrams additionally show the mean of empirical errors. This accounts for the fact that, when only the standard deviation of errors is shown, a systematical shift of the error distribution is not visible.
Beside explicitly analyzing calibration, uncertainty estimations can also be evaluated by their performance when utilized in AL schemes. While a wide variety of AL schemes exist in the MLIP literature, in this work we restrict ourselves to a simple and widely used AL approach from standard machine learning literature, known as uncertainty sampling [44, 9]. Uncertainty sampling samples a fixed configuration space, represented in a fixed dataset. An initial model is trained on a small random subset of the dataset, and the uncertainty estimation of this model is then used to iteratively select samples from the remaining pool. In the general ML literature, uncertainty sampling is often applied to classification problems, but it is also used for regression problems [45]. It was shown that uncertainty sampling can improve data efficiency, which refers to the reduction in the number of training samples needed to achieve comparable model accuracy compared to using randomly selected samples, by a multiple. However, the extent of this improvement strongly depends on the dataset and the uncertainty measure used [45]. In this study, uncertainty sampling is applied to analyze the benefit of the discussed different uncertainty measures on AL.
When it comes to GPR-based MLIPs, the calibration of uncertainty estimations has rarely been evaluated explicitly. Aside from qualitative visualizations of calibration, such as plotting errors against uncertainties[3, 46], Tran et al. [40] evaluated different aspects of uncertainty estimations for an exemplary dataset using, among others, a GPR-based MLIP. More often, uncertainty estimations of GPR-based MLIPs are evaluated by directly examining the performance when applied to AL, without explicitly analyzing calibration[3, 9, 1, 16]. However, these studies give a mixed picture on the effect of AL on training GPR-based MLIPs, and it is not always clear whether AL strategies provide an actual improvement over random sampling.
The objective of this study is to provide a comprehensive analysis of uncertainty estimations for GPR-based MLIPs. To achieve this, we provide an explicit evaluation of calibration and investigate the performance of different uncertainty measures applied to active learning (AL). We analyze global calibration using calibration curves and local calibration of groups of predictions, separated by their uncertainty, with extended reliability diagrams. Uncertainty sampling is used as a simple AL algorithm. We consider a GPR model with global Coulomb representations[22] as inputs, as well as an atomistic GPR model with SOAP[25] representations. Our tests include predicting ground-state energies for molecules from well-established benchmarks. Additionally, we predict excited-state energies for a trajectory of porphyrin in the gas phase, chosen as a large and complex molecule with a potentially complex configuration space.
@@ 2.1 Gaussian process regression
In GPR[34], a prior probability distribution over functions is defined using a Gaussian process. We want to find a posterior predictive distribution from training inputs $\mathbf{x}_{i}$ and corresponding output measurements $y_{i}$, with which we can predict outputs for unknown inputs x. Our output measurements will always be noisy. This can be considered by assuming that the actual function values differ from the measurements with
@@
where $\epsilon_{i}$ are independent, identically distributed random variables drawn from a Gaussian distribution with zero mean and variance $\sigma_{n}^{2}$. The GPR framework provides a predictive distribution for the output corresponding to any input x. The actually predicted value for the output of an input $\mathbf{x}$ is the mean of this predictive distribution, given as
@@
which is one of the most common choices. We used a different length-scale for every feature, known as automatic relevance determination[34], to increase the accuracy and flexibility of the model. Here $\mathbf{M}$ is a diagonal matrix $\text{diag}\left(\frac{1}{l_{1}^{2}},...,\frac{1}{l_{D}^{2}}\right)$, with $D$ the number of features, and $\sigma_{f}^{2}$ is the output-scale. We optimized the hyperparameters $\sigma_{f}^{2}$ and $\sigma_{n}^{2}$, as well as length-scales inside the covariance function via maximum marginal log-likelihood, details can be found in the section hyperparameter optimization of the supplementary material. For all our tests, we used the GPR implementation of GPyTorch [47].
In the case of a global representation as inputs $\mathbf{x}$, the covariance function of Eq. 3 is directly applied to these inputs. For our study, we incorporated Coulomb matrices[22] as a global representation. For Coulomb matrices, to ensure invariance against atom indexing, often a sorting is carried out, which results in discontinuities, which may cause challenges in the construction of ML models [29]. In our tests, we ignored invariance against atom permutation and omitted sorting because, for all tasks, a system with fixed atoms in a fixed order was given.
As a local representation, we used Smooth overlap of atomic positions (SOAP) [25]. It is commonly used and GPR models with SOAP representations as inputs have shown high accuracy [29]. For atomistic representations, an input x will consist of representations $\textbf{x}_{j}$ for every atomistic environment of the system. The covariance function between two inputs x and x’ is built by summing over an atomistic covariance function
@@ 2.2 Uncertainty estimations
with a given mean $m(\mathbf{x})$, and a standard deviation $u(\mathbf{x})$, which we consider as uncertainty measure in here. Note that we may interpret the pair from eq. 5 as the parameters describing a Gaussian distribution. In fact, the predictive distribution of a GPR model is anyway Gaussian. Moreover, the predictive distributions of GPR ensembles are often assumed to be Gaussian too[43, 49], which is, according to the central limit theorem, at least true for a large number of ensemble models[49].
@@
where $m_{G}(\mathbf{x})$ is the GPR mean from eq. 2 and we have the GPR standard deviation
$u_{G}\left(\textbf{x}\right)=\sqrt{k\left(\mathbf{x},\mathbf{x}\right)-\mathbf{K}\left(\mathbf{x},\mathbf{X}\right)\bigl{(}\mathbf{K}\left(\mathbf{X},\mathbf{X}\right)+\sigma_{n}^{2}\mathbf{I}\bigr{)}^{-1}\mathbf{K}\left(\mathbf{X},\mathbf{x}\right)}\,.$ (7)
Note that the GPR variance is only calculated from the kernel function $k(\cdot,\cdot)$, the inputs $\mathbf{X}$ of the training set, i.e. the sample locations in the feature space, and the noise variance $\sigma_{n}^{2}$. Hence it does not depend on the training outputs, i.e. the trained energies. Intuitively, the GPR variance is high, when $\mathbf{x}$ is far from the training inputs, as measured by the GPR model. In the variance, see eq. (7), GPR already accounts for the noise $\sigma_{n}^{2}$ in the observations, therefore, we do not add this noise to our uncertainty estimation, like it was done by Vandermause et al.[3].
The two set approach, proposed by Uteva et al.[9], was motivated by the principles underlying the jackknife and bootstrap methods[35]. We use for its predictive distribution the notation
@@
is the uncertainty estimation. The square of this particular uncertainty measure can be interpreted as a very rough estimate of the model’s variance with respect to changes in the training set, including the energies. In other words, it very roughly approximates the variance in the bias-variance decomposition of ML models, which is common in ML literature[50].
An extension of this is the bootstrap aggregation ensemble[35], which leads to the predictive distribution
@@
${m}_{b}\left(\textbf{x}\right)=\frac{1}{N}\sum_{l=1}^{N}m^{\left(l\right)}\left(\textbf{x}\right)\,,$ (12)
@@
As a slight change towards the classical definition of the two-sets and bootstrap approach, we do not use the ensemble mean as mean for their predictive distributions, since we found whether we use the mean of the ”full” GPR model or the ensemble mean does not make a difference for the essential results on the correlation between error and uncertainty estimation. A comparison can be found in Fig. S1 and S2 of the supplementary material.